OpenStudy (anonymous):

A car travels at a constant speed around a circular track whose radius is 2.77 km. The car goes once around the track in 361 s. What is the magnitude of the centripetal acceleration of the car? A car travels at a constant speed around a circular track whose radius is 2.77 km. The car goes once around the track in 361 s. What is the magnitude of the centripetal acceleration of the car? @Mathematics

OpenStudy (anonymous):

Is there a mass to this question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Given. V= 361 s R= 2.77 km ac=? ac=v^2/r ac= 361^2/2.77 This might be wrong because i didnt convert the second or km into certain unit.

OpenStudy (anonymous):

but 361 is time and not velocity